OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, -3) and (3, -5). -1 1 0 one half

OpenStudy (kagıtucak):

m=differences between y/differences between x m=-2/2=-1

OpenStudy (anonymous):

okay thanks so much! can u help me make sure i got some others right please?

OpenStudy (anonymous):

I said B on this one A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 45 dumpster loads of waste removed. In week 10, there were 30 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week. f(x) = - 3x + 45 f(x) = 3x + 45 f(x) = - 3x + 60 f(x) = 3x + 60

OpenStudy (anonymous):

@kagıtucak

OpenStudy (kagıtucak):

it is c

OpenStudy (anonymous):

oh thank! up for more? :)

OpenStudy (kagıtucak):

because after 5 week your value decreases 15 not increasing so you must have negative slope

OpenStudy (anonymous):

I said B on this one too Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1). y + 3 = -5(x - 2) y - 2 = -5(x + 3) y + 3 = -one fifth(x - 2) y - 2 = -one fifth(x + 3) and thanks for explaining

OpenStudy (kagıtucak):

m=1-2/2+3=-1/5 and if you put (2,1) in one hem you can easily see it is D

OpenStudy (anonymous):

okay just like 3 or 4 more to make sure there right if you dont mind

OpenStudy (anonymous):

I said A on this one but A and D look right Which of the following will give you the incorrect slope? the quantity y subscript two minus y subscript one over the quantity x subscript two minus x subscript one the quantity y subscript two minus y subscript one over the quantity x subscript one minus x subscript two the quantity y subscript one minus y subscript two over the quantity x subscript one minus x subscript two rise over run

OpenStudy (anonymous):

@kagıtucak